### v2

# Anisotropy Scaling Functions in Heavy-Ion Collisions: Insights into the ‘Ultra-Central Flow Puzzle’ and Constraints on Transport Coefficients and Nuclear Deformation

Abstract
Anisotropy scaling functions derived from comprehensive measurements of transverse momentum- and centrality-dependent anisotropy coefficients $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ in Pb+Pb collisions at 5.02 and 2.76 TeV, and Xe+Xe collisions at 5.44 TeV at the LHC, offer new insights into the ‘ultra-central flow puzzle’. These functions integrate diverse measurements into a single curve, clarifying anisotropy attenuation throughout the entire $p_{T}$ and centrality range. They reveal the influence of initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}$), radial flow, viscous correction to the thermal distribution function ($\delta_{f}$), the medium’s stopping power ($\hat{q}$), and specific shear viscosity ($\eta/s$) on the observed anisotropies. This analysis not only enhances understanding of transport coefficients but also provides crucial constraints on nuclear deformation.
pacs:
25.75.-q, 25.75.Dw, 25.75.Ld
Azimuthal anisotropy measurements are pivotal in the study of the quark-gluon plasma (QGP) produced in heavy ion collisions at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC). These measurements, crucial for determining the temperature ($T$) and baryon chemical potential ($\mu_{B}$) dependence of transport coefficients, are quantified by the complex coefficients Bilandzic et al. (2011); Luzum (2011); Teaney and Yan (2012):
$V_{n}\equiv v_{n}e^{in\Psi_{n}}=\langle e^{in\phi}\rangle,$ (1)
[…]
These $v_{n}$ values are linked to the Fourier coefficients $v_{nn}$, which describe the intensity of two-particle correlations in azimuthal angle differences $\Delta\phi=\phi_{\mathrm{i}}-\phi_{\mathrm{j}}$ Poskanzer and Voloshin (1998); Lacey (2006):
$\frac{dN^{pairs}}{d\Delta\phi}\propto 1+2\sum_{n=1}^{\infty}v_{nn}\cos(n\Delta\phi),$
$v_{nn}(p_{T}^{i},p_{T}^{j})=v_{n}(p_{T}^{i})v_{n}(p_{T}^{j})+\delta_{\text{NF}},$ (2)
where $\delta_{\text{NF}}$ indicates non-flow contributions, minimized through specific experimental methodologies Lacey (2006); Luzum and Ollitrault (2011); Retinskaya et al. (2012); Aad et al. (2012).
The coefficients $v_{n}$ are fundamentally linked to collective flow dynamics for $p_{T}\lesssim 4-5$ GeV, transitioning to jet quenching at higher $p_{T}$. This behavior is supported by extensive research Song et al. (2011); Alver et al. (2008, 2010); Ollitrault et al. (2009); Dusling et al. (2010); Lacey et al. (2010); Shen et al. (2011); Niemi et al. (2013); Fu (2015); Andrés et al. (2015); Abdallah et al. (2022); Adam et al. (2018, 2016a); Qiu and Heinz (2011); Adare et al. (2011); Magdy (2019); Adamczyk et al. (2016a, b, 2015, c); Adam et al. (2019); Gardim et al. (2015); Holopainen et al. (2011); Qin et al. (2010); Qiu and Heinz (2011); Gale et al. (2013); Burke et al. (2014); Liu and Lacey (2018); Adcox et al. (2002); Adler et al. (2002); Zhang et al. (2008); Abelev et al. (2014); Mehtar-Tani et al. (2013); Qin and Wang (2015), which highlights the impact of radial expansion, $v_{n}$-fluctuations, $p_{T}$-dependent viscous attenuation, jet quenching, and initial-state anisotropy on $v_{n}$. The transverse plane density profile $\rho(r,\varphi)$ is quantified by complex eccentricity coefficients:
${\mathcal{E}_{n}\equiv\varepsilon_{n}e^{in\Phi_{n}}=\frac{\int d^{2}r_{\perp}\,r^{m}\,e^{in\varphi}\,\rho(r,\varphi)}{\int d^{2}r_{\perp}\,r^{m}\,\rho(r,\varphi)}},$ (3)
where $r$ and $\varphi$ represent the radius and azimuthal angle, respectively, and ${\Phi_{n}}$ denotes the angle of the ${n^{\rm th}}$-order participant plane (${m=n}$ for ${n{\geq}2}$ and ${m=3}$ for ${n=1}$) Qiu and Heinz (2011); Fu (2015); Niemi et al. (2016); Noronha-Hostler et al. (2016). The deformed Woods-Saxon distribution, employed to model nucleon configurations in non-spherical nuclei, is described by:
$\rho(r,\theta,\varphi)$ $=\frac{\rho_{0}}{1+\exp\left(\frac{r-R(\theta,\varphi)}{a}\right)},$ (4)
$R(\theta,\varphi)$ $=R_{0}\left\{1+\beta_{2}\left[\cos\gamma Y_{20}(\theta,\varphi)+\sin\gamma Y_{22}(\theta,\varphi)\right]\right\},$
where $\rho_{0}$ is the central density, $R_{0}$ the nuclear radius, and $a$ the skin thickness. The spherical harmonics $Y_{lm}$, along with coefficients $\beta_{2}$ and $\gamma$, shape the nucleus, with $\beta_{2}$ measuring deformation magnitude and $\gamma$ indicating asymmetry degrees between 0 and $60^{\circ}$. Fluctuations in the initial state density profile lead to variations in $\varepsilon_{n}$.
Dynamical models based on relativistic hydrodynamics, which suggest a roughly linear relationship $v_{n}\propto\varepsilon_{n}$ Noronha-Hostler et al. (2016); Sievert and Noronha-Hostler (2019); Rao et al. (2021), have successfully replicated the observed magnitudes and trends of $v_{n}$ coefficients Qin et al. (2010); Schenke et al. (2011); Qiu and Heinz (2011); Shen et al. (2011); Bozek (2012); Gardim et al. (2012); Hirano et al. (2013). These models are also central to efforts employing Bayesian inference to constrain the transport properties of the QGP Bernhard et al. (2016, 2019); Moreland et al. (2020); Everett et al. (2021a, b); Nijs et al. (2021); Auvinen et al. (2020); Parkkila et al. (2021). However, they encounter significant challenges in ultra-central collisions (cent $\lesssim 1\%$), where they struggle to simultaneously predict $v_{2}$ and $v_{3}$ accurately Giannini et al. (2023). Typically, these models either overestimate $v_{2}$ or underestimate $v_{3}$, or display both discrepancies.
The inability of hydrodynamic models to accurately describe both $v_{2}$ and $v_{3}$ in ultra-central collisions presents a significant challenge, contradicting the expectations set for these collisions. Conventionally, models were expected to perform well in central collisions due to higher charge particle multiplicities ${\left\langle{{\text{N}}_{\text{chg}}}\right\rangle}$ and larger volumes of locally thermalized domains. This discrepancy raises doubts about whether these widely used models are missing crucial components necessary for accurately depicting the initial state and transport coefficients.
Despite extensive efforts to resolve the ‘ultra-central flow puzzle’ by refining initial conditions Luzum and Ollitrault (2013); Denicol et al. (2014); Shen et al. (2015); Bhalerao et al. (2015); Loizides (2016); Giacalone et al. (2019); Gelis et al. (2019); Carzon et al. (2020); Snyder et al. (2021); Zakharov (2020), examining transport coefficients Luzum and Ollitrault (2013); Rose et al. (2014); Shen et al. (2015); Plumari et al. (2015), and evaluating equations of state Alba et al. (2018), a definitive resolution remains elusive. The ongoing investigations highlight the complexity of accurately modeling these critical aspects of QGP behavior.
In this study, we develop anisotropy scaling functions, premised on the idea that diverse measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ can be unified into a single, coherent scaling function. This unification integrates various parameters affecting $v_{n}(p_{T},\text{cent})$, including initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}\propto RT$), radial flow magnitude, the medium’s stopping power ($\hat{q}$), and the specific shear viscosity or viscosity-to-entropy ratio ($\eta/s\propto T^{3}/\hat{q}$), along with the viscous correction to the thermal distribution function ($\delta_{f}$ajumder et al. (2007); Dusling et al. (2010).
These parameters contribute to the expression of the anisotropy coefficients, which can be described by Eq. 5 Staig and Shuryak (2011); Gubser and Yarom (2011); Lacey et al. (2013); Liu and Lacey (2018):
$v_{n}(p_{T},\mathrm{cent})=\varepsilon_{n}e^{-\frac{\beta}{\mathbb{R}}\left[n(n+\kappa p_{T}^{2})\right]},\,\,n=2,3,$ (5)
where $\beta\propto\eta/s$, $\delta_{f}=\kappa p_{T}^{2}$ Dusling et al. (2010); Liu and Lacey (2018), and $\mathbb{R}\propto\left\langle{{\text{N}}_{\text{chg}}}\right\rangle_{|\eta|\leq 0.5}^{1/3}$ is related to the mid-rapidity (${|\eta|\leq 0.5}$) charged particle multiplicity. The transition from flow to jet quenching at higher $p_{T}$ is managed by maintaining consistency between $\eta/s$ and $\hat{q}$, ensuring a smooth transition in the low and high momentum regions of $\delta_{f}$. This shift is facilitated by fixing the $\kappa p_{T}^{2}$ term in Eq. 5 to remain constant for $p_{T}$ values above approximately $4.5$ GeV/c, marking the threshold between flow-dominated and jet-quenching domains.
Equation 5 reveals a crucial insight: In the domain of the most central collisions, characterized by the peak value $\mathbb{R}_{0}$, a scaling connection emerges. This equation links the observed values of the harmonic $v_{n}(p_{T},0)$ in these ultra-central events with those observed as $v^{{}^{\prime}}_{n}(p_{T},\text{cent})$ at different centralities, each tied to its corresponding $\mathbb{R}^{{}^{\prime}}$ value:
$\frac{{v_{n}}(p_{T},\mathrm{0})}{{{\varepsilon}_{n}(0)}}{{e}^{+\frac{n\beta}{\mathbb{R}_{0}}(n+\kappa p_{T}^{2})}}=\left(\frac{v_{n}^{{}^{\prime}}({p_{T}},\mathrm{cent})}{\varepsilon_{n}^{{}^{\prime}}(\text{cent})}\right){{e}^{+\frac{n\beta}{\mathbb{R}_{0}}(n+\kappa p_{T}^{2})\left(\frac{{{\mathbb{R}_{0}}}}{{{\mathbb{R}}^{{}^{\prime}}}}-1\right)}},$ (6)
where the exponential on the left-hand side of the equation accounts for the anisotropy attenuation in ultra-central collisions, and that on the right-hand side accounts for the relative attenuation at other centralities.
Figure 1: (Color Online) Panel (a) contrasts the centrality-dependent values of $\varepsilon_{2}(\text{cent})$ and $\varepsilon_{3}(\text{cent})$ for 5.02 TeV Pb+Pb collisions featuring spherical Pb nuclei and 5.44 TeV Xe+Xe collisions for deformed Xe nuclei. In panel (b), the ratios $\varepsilon_{3}(\text{cent})/\varepsilon_{2}(\text{cent})$ are compared for spherical Pb nuclei ($\beta_{2}=0$) and both spherical and deformed Xe nuclei (cf. Eq. 4), as indicated.
Moreover, when preserving a fixed centrality within a given system, a revealing correlation emerges between the values of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$. These values, both subject to the same influences of radial flow, $\eta/s$, and $\hat{q}$, adhere to the following relationship:
$\frac{{v_{2}}({p_{T}},\text{cent})}{{\varepsilon_{2}}(\text{cent})}{e^{\frac{2\alpha\beta}{\mathbb{R}_{0}}}}={{\left(\frac{v_{3}({p_{T}},\text{cent})}{\varepsilon_{3}(\text{cent})}\right)}^{\frac{2}{3}}},$ (7)
where $\alpha$ represents a system-dependent (yet centrality-independent) normalizing constant. Equations 6 and 7 encapsulate the intricate dependencies of the high and low $p_{T}$ anisotropy coefficients from ultra-central to peripheral collisions, underlining that the scaled values of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ measurements can be effectively collapsed onto a single curve or scaling function $S\!{{}_{\text{FS}}}$. Therefore, the identification of such a scaling function would furnish compelling evidence for the coherence of the scaling coefficients and the trustworthiness of the corresponding eccentricity spectrum and its ratios.
Figure 2: (Color Online) Comparison of $v_{2}(p_{T})$ and $v_{3}(p_{T})$ in panel (a), their eccentricity-scaled values $v_{2}/\varepsilon_{2}$ and $v_{3}/\varepsilon_{3}$ in panel (b), and the resulting scaling function in panel (c) for 0.1% central Pb+Pb collisions at 5.02 TeV. Panel (c) also includes scaled results for 30-40% central Pb+Pb collisions. The data are sourced from the ATLAS collaboration Aaboud et al. (2018).
The eccentricity ratio, as shown in Fig. 1, is crucial for understanding the nuanced impact of centrality and deformation on collision dynamics. This figure displays the calculated values of $\varepsilon_{2}(\text{cent})$, $\varepsilon_{3}(\text{cent})$, and their ratios for Pb+Pb and Xe+Xe collisions at 5.02 TeV and 5.44 TeV. These values reveal distinct dependencies that not only help refine the eccentricity spectrum but also illuminate the initial-state deformation of the Xe nucleus. Particularly in ultra-central and central collisions, these measurements are vital for accurately quantifying nuclear deformation. Additionally, the ratio $\varepsilon_{2}(\text{cent})/\varepsilon_{3}(\text{cent})$ demonstrates insensitivity to variations in $\gamma$, highlighting its stability across different asymmetries.
Figure 3: (Color Online) Comparison of $v_{2}(p_{T},\rm{cent})$ in panel (a), their eccentricity-scaled values [$v_{2}(p_{T},\rm{cent})/\varepsilon_{2}(\text{cent})$] in panel (b), and the resulting scaling function in panel (c) for Pb+Pb collisions at 5.02 TeV. Data sourced from the ALICE collaboration Acharya et al. (2018a).
Figure 4: (Color Online) Comparison of the $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ values for Xe+Xe collisions and $v_{2}(p_{T},\text{cent})$ for Pb+Pb collisions in panel (a), their eccentricity-scaled values $v_{2}(p_{T},\text{cent})/\varepsilon_{2}(\text{cent})$ and $v_{3}(p_{T},\text{cent})/\varepsilon_{3}(\text{cent})$ in panel (b), and the resulting scaling function in panel (c) for 0-5% central collisions. Panel (c) also includes scaled results for 30-40% central Xe+Xe collisions.The data are sourced from the ALICE collaboration Acharya et al. (2018b, a) .
The data utilized in this study are sourced from the ATLAS Aad et al. (2012); Aaboud et al. (2018) and ALICE Acharya et al. (2018b, a) collaborations, encompassing $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ measurements for Pb+Pb collisions at $\sqrt{s_{NN}}=2.76$ and $5.02$ TeV, as well as Xe+Xe collisions at $5.44$ TeV. The required centrality-dependent $\left\langle{{\text{N}}_{\text{chg}}}\right\rangle_{|\eta|\leq 0.5}$ values were derived from corresponding multiplicity density measurements Aamodt et al. (2011); Adam et al. (2016b); Acharya et al. (2019); Sirunyan et al. (2019). We adopted the previously established value $\kappa=0.17$ (GeV/c)-2 Liu and Lacey (2018) to compute $\delta{f}$. The eccentricities were computed according to the procedure outlined in Eq. 3, with the assistance of a Monte Carlo quark-Glauber model (MC-qGlauber) featuring fluctuating initial conditions Liu and Lacey (2018). This model, based on the widely used MC-Glauber model Miller et al. (2007); Alver et al. (2007), accounts for the finite size of the nucleon, the nucleon’s wounding profile, the distribution of quarks within the nucleon, and quark cross sections that accurately reproduce the NN inelastic cross section for the corresponding beam energies. Calculations were performed for spherical Pb nuclei and Xe nuclei with varying degrees of initial-state deformation characterized with different values for the $\beta_{2}$ and $\gamma$ parameters (cf. Eq. 4). A systematic uncertainty of 2-3% was estimated for the eccentricities based on variations in the model parameters.
The scaling functions were derived from differential measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ across a range of collision energies, including Pb+Pb collisions at 2.76 and 5.02 TeV, and Xe+Xe collisions at 5.44 TeV. This derivation utilized specific equations (6 and 7), with the reference value $\mathbb{R}_{0}$ for ultra-central Pb+Pb collisions at 5.02 TeV. In Figure 2, we illustrate the scaling procedure for 0.1% central Pb+Pb collisions at 5.02 TeV, employing $1/\sqrt{p_{T}}$ for the x-axis in panels (b) and (c) to highlight the flow- and jet-quenching-dominated domains. Panel (a) emphasizes the discrepancy between $v_{2}(p_{T})$ and $v_{3}(p_{T})$, while panel (b) shows that eccentricity scaling alone ($v_{2}(p_{T})/\varepsilon_{2}$ and $v_{3}(p_{T})/\varepsilon_{3}$) fails to capture their difference. Panel (c) showcases the resulting scaling function, indicating convergence of data onto a single curve for both $p_{T}$ domains, providing robust evidence for the consistency of the scaling coefficients and the validity of the corresponding eccentricity spectrum and its ratios. Validation of the scaling function spanned the entire measurement range (0.0-0.1% to 50-60%) for specific parameters ($\beta=0.88$ and $\alpha=1$), as demonstrated in Figure 3 for $v_{2}(p_{T},\text{cent})$ measurements. The attenuation factor for the 0-5% centrality cut reflects an average incorporating values from both ultra-central and non-ultra-central collisions, considering contributions from both scenarios.
A similarly robust scaling function, closely matching that for Pb+Pb collisions at 5.02 TeV, was obtained across the full range of measurements for Pb+Pb collisions at 2.76 TeV. Characterized by scaling exponents $\beta=0.84$ and $\alpha=1.25$, the results indicate an approximate 5% reduction in $\eta/s$ when the collision energy decreases from 5.02 TeV to 2.76 TeV. Importantly, the consistency of the scaling functions at both energies suggests a negligible influence from non-flow effects.
The observed scaling function, consistent across both flow and jet-quenching domains with identical scaling coefficients, not only constrains $\eta/s$ and $\hat{q}$ but also supports the relationship between $\eta/s$ and $T^{3}/\hat{q}$ proposed by Majumder et al. Majumder et al. (2007). This scaling, notably the high $p_{T}$ anisotropy’s dependence on the dimensionless size $\mathbb{R}\propto\left\langle{\text{N}_{\text{chg}}}\right\rangle^{1/3}$, underscores the influence of path length and highlights the role of radiative energy loss in shaping jet-quenching-induced anisotropy. Moreover, the seamless transition observed between the low and high momentum regions at the $p_{T}$ threshold, which delineates flow from jet-quenching domains, provides crucial insights into the relative magnitude of $\eta/s$ and $T^{3}/\hat{q}$ Dusling et al. (2010). This distinction is critical for identifying the nature of the quark-gluon plasma as either strongly or weakly coupled. A preliminary analysis, following the $\delta f$ formalism in Ref. Dusling et al. (2010) and reconciling $\eta/s$ with $\hat{q}$, suggests that $\eta/s$ exceeds $T^{3}/\hat{q}$, indicating a strongly coupled plasma.
The scaling function elucidates nuclear deformation, as illustrated in Figs. 1 and 4. In Fig. 4, we compare $v_{2}(p_{T})$ and $v_{3}(p_{T})$ from Xe+Xe collisions against $v_{2}(p_{T})$ from Pb+Pb collisions at 0-5% centrality, using eccentricities for the deformed Xe nucleus ($\beta_{2}=0.16$ Tsukada et al. (2017)). Panel (a) highlights significant differences in viscous attenuation between these harmonics, with Xe+Xe showing larger magnitudes, aligned with its higher eccentricities (cf. Fig. 1). Panel (b) acknowledges that eccentricity scaling does not entirely account for these differences but clarifies the contrast in $v_{2}(p_{T})$ between Xe and Pb. Panel (c) demonstrates the scaling function’s ability to merge data into a single curve across both flow- and jet-quenching-dominated regimes, validating the consistency of scaling coefficients and the accuracy of the eccentricity spectrum for the deformed Xe nucleus. This spans centrality ranges from 0-5% to 40-50%, with $\beta=0.90$, $\alpha=0$, and $\beta_{2}=0.16$. Deviations from these values suggest mild deformations in Xe nuclei, correlating with an approximate 2% increase in $\eta/s$ when increasing the collision energy from 5.02 TeV to 5.44 TeV. Further analysis shows slightly higher radial flow in 5.02 TeV Pb+Pb collisions compared to 5.44 TeV Xe+Xe collisions.
The application of scaling functions to analyze nuclear deformation significantly enhances the precision of our studies. This method (i) compensates effectively for anisotropy attenuation in ultra-central collisions, (ii) decouples initial- and final-state effects, and (iii) relies on consistent measurements of $v_{2}(p_{T})$ and $v_{3}(p_{T})$ across events with identical $\eta/s$, multiplicity, and radial flow characteristics. Consequently, it markedly reduces systematic uncertainties.
In summary, anisotropy scaling functions, derived from extensive measurements of $v_{2}(p_{T},\text{cent})$ and $v_{3}(p_{T},\text{cent})$ in Pb+Pb collisions at 5.02 and 2.76 TeV, and Xe+Xe collisions at 5.44 TeV, provide crucial insights into the ‘ultra-central flow puzzle’ and constrain transport coefficients, the eccentricity spectrum, and nucleus deformation. Consolidating diverse measurements into a single curve, these functions account for specific parameters influencing $v_{n}(p_{T},\text{cent})$, including initial-state eccentricities ($\varepsilon_{n}$), dimensionless size ($\mathbb{R}$), radial flow magnitude, the medium’s stopping power ($\hat{q}$), and its specific shear viscosity ($\eta/s$), as well as the viscous correction to the thermal distribution function ($\delta_{f}$). They yield distinct constraints for transport coefficients and nucleus deformation.
Acknowledgement
This research is supported by the US DOE under contract DE-FG02-87ER40331.A008.